OpenStudy (anonymous):

Find the roots of the polynomial equation. x^3-x^2+x+39 1+3i, 1-3i, -3 1+3i, 1-3i, 3 -2+3i, -2-3i, -3 2+3i, 2-3i, -3

OpenStudy (solomonzelman):

it will be +- factors of 39, plug each in to see which of them work.

OpenStudy (anonymous):

the factors would be 1, 3, 13, 39 right?

OpenStudy (solomonzelman):

yes, the factors, yes, but the roots can be negative or positive.

OpenStudy (mathstudent55):

Plug in x = 3 and x = -3 into the polynomial. Whichever of the two gives you zero is a root. Call it k. Then using long division or synthetic division, divide the original polynomial by x - k to get a quadratic equation. Then use the quadratic formula to solve the quadratic equation.

OpenStudy (anonymous):

I plugged in 3 and -3 both don't equal 0 @mathstudent55

OpenStudy (mathstudent55):

Try again. One of the two definitely makes it zero.

OpenStudy (anonymous):

One is 60 and the other is 18

OpenStudy (mathstudent55):

\(x^3-x^2+x+39 = 0\) \( (3)^3 - (3)^2 + (3) + 39 \) \( (-3)^3 - (-3)^2 + (-3) + 39 \)

OpenStudy (mathstudent55):

Let's do each one. (3)^3 - (3)^2 + (3) + 39 = 27 - 9 + 3 + 39 = 60 Your 60 is correct.

OpenStudy (mathstudent55):

(-3)^3 - (-3)^2 + (-3) + 39 = -27 - 9 - 3 + 39 = 0 Your 18 is incorrect. Remember that (-3)^3 = -27 and -(-3)^2 = - 9

OpenStudy (anonymous):

Oh, now I divide the equation by x-3?

OpenStudy (mathstudent55):

Be careful. When a polynomial equation has roots, a, b, c, d, etc. The binomial factors are (x - a), (x - b), (x - c), (x - d), etc. Since the root is -3, you need x - (-3) = x + 3. Divide the polynomial by x + 3.

OpenStudy (anonymous):

Oh I see.

OpenStudy (anonymous):

After synthetic division I got 1 4 13 0

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

I get \(x^2 - 4x + 13\) with long division. Can you check on your synthetic division?